The New England Merchants Bank Building in Boston is 152 m high. On windy days it sways with a frequency of 0.17 Hz , and the acceleration of the top of the building can reach 1.7 % of the free-fall acceleration, enough to cause discomfort for occupants. What is the total distance, side to side, that the top of the building moves during such an oscillation?

(2A= ??? )
Answer #1

As we know that;

= f=0.17 Hz

Then

angular frequency is ,

= w=2πf= 2*π*0.17

= w= 1.068 rad/sec

Then,

= amax= 1.7% of g

= 0.017*9.8 = 0.167 m/s^2

Then

Assume A amplitude of vibrations,

= amax= Aw^2

= A= amax/w^2

= A= 0.167/1.068^2

= A= 0.146 m

Then,

The total distance ,side to side = 2A

=2*0.146

= 0.292 m

= 29.2 cm